# Microcanonical simulated annealing: Massively parallel Monte Carlo simulations with sporadic random-number generation

### v1

@@ Abstract
Numerical simulations of models and theories that describe complex experimental systems —in fields like high-energy and condensed-matter physics— are becoming increasingly important. Examples include lattice gauge theories, which can describe, among others, quantum chromodynamics (the Standard Model description of strong interactions between elementary particles), and spin-glass systems. Beyond fundamental research, these computational methods also find practical applications, among many others, in optimization, finance, and complex biological problems. However, Monte Carlo simulations, an important subcategory of these methods, are plagued by a major drawback: they are extremely greedy for (pseudo) random numbers. The total fraction of computer time dedicated to random-number generation increases as the hardware grows more sophisticated, and can get prohibitive for special-purpose computing platforms. We propose here a general-purpose microcanonical simulated annealing (mic.SA) formalism that dramatically reduces such a burden. The algorithm is fully adapted to a massively parallel computation, as we show in the particularly demanding benchmark of the three-dimensional Ising spin glass. We carry out very stringent numerical tests of the new algorithm by comparing our results, obtained on GPUs, with high-precision standard (i.e., random-number-greedy) simulations performed on the Janus II custom-built supercomputer. In those cases where thermal equilibrium is reachable (i.e., in the paramagnetic phase), both simulations reach compatible values. More significantly, barring short-time corrections, a simple time rescaling suffices to map the mic.SA off-equilibrium dynamics onto the results obtained with standard simulations.
@@ I Introduction
Here we propose a widely usable microcanical Simulated Annealing (mic.SA) formalism, which is both extremely frugal in the number of needed random numbers and fully adapted to a massively parallel computation. The method is demonstrated in the rather demanding case of the three-dimensional Ising spin glass. We carry out our simulations on GPUs to demonstrate that our algorithm nicely parallelizes. We start from a microcanonical ensemble for an extended configuration space that comprises the degrees of freedom we are really interested in (the spins) and some auxiliary degrees of freedom traditionally called daemons. We shall term these auxiliary variables walkers if they are allowed to wander through the system or daemons if they are kept tied to a single lattice site; should the distinction be immaterial, which happens at many points in the discussion, we shall name them daemons/walkers. The reader will find useful analogies between our Monte Carlo formalism and molecular dynamics simulations coupled with a thermostat [30].
@@
Figure 1: After a time rescaling, our microcanonical simulated annealing reproduces canonical dynamics. We study the energy relaxation towards its equilibrium value by plotting the excess energy per spin over its equilibrium value, $e(t)-e_{\infty}$, versus simulation time $t$, as computed for a spin-glass sample on a $160\times 160\times 160$ lattice with binary couplings at temperature $T=0.7\approx 0.64T_{\mathrm{c}}$ ($T_{\mathrm{c}}$ is the critical temperature for the phase transition separating the paramagnetic phase at high temperatures from the spin-glass phase at low temperature). The sample is investigated using two dynamics: a standard, random-number greedy Metropolis algorithm (the label can in the plot stands for canonical; these data were obtained from the Janus II supercomputer) and the new microcanonical simulated annealing (mic.SA) algorithm introduced in this work. Specifically, we are using the version of the algorithm with 6 walkers per spin, see Sec. II.8. We plot both data sets, subtracting the same estimate for $e_{\infty}$, which was obtained from the canonical simulation as explained in Appendix A. After the mic.SA time is rescaled by the factor $k^{*}=6.034$ (see Table 1), $t^{\text{plot}}=k^{*}\times t^{\text{mic.SA}}$, both energy relaxations are indistinguishable (the naif rescaling would have been $k^{\text{naif}}=6$, because of the 6 walkers per spin).
@@
A microcanonical framework with just one daemon was proposed by Creutz [31], precisely to minimize the need for random numbers. Unfortunately, Creutz’s ensemble leads to a rather significant dynamic slowing down in the context of spin glasses [32]. More promising is Lustig’s microcanonical ensemble [33], which contains a number of daemons proportional to the number of spins. Lustig’s framework is very effective for studying first-order phase transitions [34], also for disordered systems [35], which suggests a potential effectiveness in the study of spin glasses. This formalism, however, has three major inconveniences: (i) it is unsuitable for parallel computations, (ii) it is not frugal in its usage of random numbers and (iii) it eventually approaches the equilibrium state corresponding to the microcanonical ensemble (which coincides with the canonical ensemble only in the limit of an infinite system size). We cure problems (i) and (ii) by simply skipping the integration of the daemons that Lustig carries out and keeping them instead as explicit dynamic variables. In order to cure problem (iii), we generalize the simulated annealing algorithm [17] to the microcanonical context. We argue (and show by example) that it is sufficient to lower the total energy of the system by acting solely on the daemons at a logarithmic time rate (this is the only part of the algorithm that involves random numbers) to obtain results indistinguishable (after a simple time rescaling) from those produced using a canonical simulation employing the Janus II special-purpose computer [15]. Although rescalability will be considered in much deeper detail below, as a starter, the reader may wish to check Fig. 1.
@@ II.5 Daemons/walkers with a limited capacity
In other words, we should reject any spin-flip that lowers the energy by an amount larger than what the daemon/walker responsible for that update can currently accommodate. On the other hand, having daemons/walkers representable with just two bits reduces the total memory needed. Furthermore, in a multispin coding simulation, the whole algorithm should be implemented employing solely Boolean operations [this includes checking condition Eq. (11) or performing the update in Eq. (9)], which is certainly simpler if the daemons/walkers are representable with only two bits.
@@ II.7 The microcanonical simulated annealing
From the above discussion, it should be clear that refreshing the daemons/walkers configuration at regular times, say every 1000 timesteps (see Sec. II.8 for the precise definition of timestep), should eventually lead any system of finite size to canonical equilibrium. This would dilute the need for random numbers by a factor of 1000, which would probably make the computational weight of generating random numbers irrelevant for most (if not all) applications. However, we shall try to be even thriftier with our random numbers. Indeed, the energy relaxation of spin glasses at and below their critical temperature decreases very slowly with time $t$. The excess energy over the equilibrium value decays as a power law, $H^{\text{\tiny EA}}(t)-\langle H^{\text{\tiny EA}}\rangle_{\beta}\propto N/t^{\epsilon(T)}$, where $\epsilon(T)$ is a fairly small exponent [55, 43] (the exponent is so small that the power law resembles a logarithmic behavior). In fact, for a law $1/\log(t)$, the time required to have a prefixed energy factor decay $f$ grows exponentially with $1/f$. We decided accordingly to refresh our daemons/walkers with only an exponential time schedule, at timesteps $t_{s}=2^{s}$ with $s$ a positive integer or zero (i.e., at timesteps $1,2,2^{2},2^{3},2^{4},2^{5},\ldots$). At each of those timesteps, the spin system will experience an abrupt, but tiny, energy change as a reaction to the cooler daemons/walkers. This is inappreciable at the scale of Fig. 1, but can be observed with the more refined analysis in Fig. 2-d.
@@ III Results
Figure 2: Comparison of Metropolis and mic.SA dynamics with 6 walkers in the spin-glass phase at $T\!=\!0.7\approx 0.64T_{\mathrm{c}}$. a) Time evolution of the difference of the energy at time $t$ and the equilibrium value obtained through a fit to the canonical data. The solid yellow line is for the best fit described in Table 2. b) The coherence length, $\xi_{12}$, as a function of the simulation time for both models. The solid line is again for the best fit described in Table 2. c) Coherence length, $\xi_{12}$, as a function of the ratio between the coherence lengths $\xi_{23}$ and $\xi_{12}$. d) Difference between the results from the two algorithms for the energy, Eq. (22), as a function of the simulation time. e) As in d) but for the coherence length. To compute these differences, we have considered the best fit to our canonical data —solid lines in a) and b)— and the rescaling coefficient $k^{*}=6.034$ (see Appendix A for more information on the estimation of $k^{*}$). Error bars are smaller than data points. For the sake of clarity points are plotted at regular time intervals in logarithmic scale.
@@
Our estimates for $e(t)$, $\xi_{12}(t)$ and the dimensionless ratio $\xi_{23}/\xi_{12}(t)$ are displayed as a function of time in Fig. 2 ($T=0.7\approx 0.64T_{\mathrm{c}}$), Fig. 3 ($T=1.1\approx T_{\mathrm{c}}$) and Fig. 4 ($T=1.4\approx 1.27T_{\mathrm{c}}$). For all three temperatures, we observe that, when $\xi_{12}(t)$ is plotted parametrically as a function of $\xi_{23}/\xi_{12}(t)$, the canonical and the mic.SA data fall onto a single curve. This finding indicates that it suffices to consider $\xi_{12}(t)$ when trying to find the correspondence between the $C_{4}(r,t)$ computed with the two dynamics. The figures suggest as well that a shift in $\log t$ (i.e., a simple rescaling for $t$) suffices to superimpose the mic.SA data with the canonical curve for both $e(t)$ and $\xi_{12}(t)$. We shall attempt to make this observation quantitatively precise.
@@
Having promoted the canonical $\xi_{12}(t,T)$ to a continuous function, we can simply fit the mic.SA curves to
$\xi^{\text{mic.SA}}_{12}(t,T)=F^{\xi}_{\text{can}}(k\,t,T)\,,$ (21)
with the rescaling factor $k$ as the only fitting parameter. The best-fit rescaling parameter is named $k^{*}$ in Table 1 and in Figs. 23, and 4. In the figures, we have used for the energy the same value of $k^{*}$ computed with $\xi_{12}$. Each of the different variants of the mic.SA algorithm will have its own $k^{*}$ that slightly depends on the temperature. For further details on the computation of $k^{*}$, see Fig. 5 and Appendix A.
Figure 5: Estimation of the best time shift $k^{*}$. In the main panel, the difference between the logarithm of the original measured time in the mic.SA data and the interpolating function $f(k^{*}t)$ as a function of the coherence length $\xi_{12}$. The horizontal solid line at zero indicates the interval of the fit used to estimate $k^{*}$ (see Table 1 for details). Inset: $\log(t)$ as a function of $\xi_{12}$ for both algorithms. Solid lines indicate the best fit for our data. In the case of the mic.SA data, the only fitting parameter is $k$, while the rest of the parameters $\Theta=\{a_{0},\,a_{1},\,z\}$ come from the fit to the canonical data described in Appendix A.
@@
$\delta_{e}(t,k)$ $=$ $e^{\text{mic.SA}}(t)-F^{e}_{\text{can}}(k\,t,T)\,,$ (22)
$\delta_{\xi}(t,k)$ $=$ $\xi_{12}^{\text{mic.SA}}(t)-F^{\xi}_{\text{can}}(k\,t,T)\,,$ (23)
as computed for $k=k^{*}$. These figures show that, when looked at with enough accuracy, $e^{\text{mic.SA}}(t)$ is not a smooth function of time (this is most clear for $T\leq T_{\mathrm{c}}$). Indeed, these curves have abrupt (but tiny) changes every time the walkers get refreshed. This is why we have not attempted to get the rescaling factor from the energy. These abrupt changes, as well as the difference $\delta_{e}(t)$ itself, strongly decrease as the time $t$ increases.
@@ IV Conclusions
We have proposed a new Monte Carlo algorithm, microcanonical simulated annealing (mic.SA), that is both fully adapted to massively parallel computations and extremely frugal in its use of random numbers. The method is fully general and has a certain number of adaptable features. Chief among these are the number $\gamma$ of daemons/walkers per elementary degree of freedom (spins in our case) and the choice between using daemons or walkers. Walkers have the advantage of avoiding local conservation laws and their resulting slower dynamics, but one can decide whether to use daemons or walkers, and how many, on the view of the application at hand and the hardware platform that will be used.
@@
The method performs fully satisfactorily in the demanding setting of the three-dimensional Ising spin glass, where we have shown that it can reproduce the results of a standard (random-number-greedy) Metropolis algorithm both in and out of equilibrium. In particular, we have shown that the out-of-equilibrium dynamics derived from mic.SA can be rescaled onto the one from traditional Metropolis by a simple change in the time unit. The fact that, in the case of walkers, the rescaling factor is nearly equal to the ratio of the number of spin updates in a single timestep for both methods (which coincides with $\gamma$) tells us that a single step of mic.SA is worth $\gamma$ Metropolis sweeps, without requiring a single random number.
@@
Our CUDA implementation demonstrates the parallel-friendly nature of the algorithm. Our code, see Appendix B, is publicly available and puts in the hands of a general user a computing performance that could be previously delivered only by the Janus II custom-built computer  [^3] In the context of quantum spin glasses, performance competitive with Janus II has been already obtained on GPUs [25].. Our program may run both standard Metropolis and mic.SA, with the mic.SA version taking half the time (in fact, in our Metropolis implementation, random-number generation takes about 75% of the simulation time). A standard metric corresponds to the number of picoseconds per spin update. Our mic.SA program runs on an H100 GPU at 0.4 ps/spin-update, versus the 3.17 ps/spin-update on a single Janus II FPGA. The above two numbers should, however, be compared with care for two reasons. First, the computing power on the GPU is spread over 128 replicas, while on Janus II updates are concentrated on only two replicas. Hence, for a high-accuracy simulation spanning a limited time range (say up to $t\sim 10^{9}$), the GPU is faster. If, however, one needs to reach times $t\sim 10^{11}$ or larger (which is not uncommon [6]), Janus II still fares better (the GPU might make more spin updates per second by running many replicas, but would need a wall-clock time of many months to finish each simulation). Second, we are comparing 2025 high-end GPUs with the 2013 Virtex 7 FPGAs that Janus II is made of, a somewhat unfair comparison.
@@
Since mic.SA is particularly suitable for high-end FPGAs or even dedicated hardware (based on an ASIC, for instance), we expect that users interested in extreme-scale Monte Carlo simulations will find a significant advantage in our proposal.
@@
The performance of mic.SA for system sizes $L$ small enough to be brought to thermal equilibrium also in the spin-glass phase, $T<T_{\mathrm{c}}$, needs to be ascertained. In particular, mic.SA needs to be combined with Parallel Tempering [61]. Although we do not foresee difficulties of principle, this algorithmic combination is yet to be tested.
The dynamical properties of mic.SA in small systems are likely to be quite peculiar. If the daemon/walkers are refreshed at a very low rate, the system may have enough time to equilibrate in the time lapse between the refreshing steps. Since the equilibrium values of the different observables differ in the canonical and the microcanonical ensemble, it is not obvious that it will always be possible to map mic.SA dynamics onto canonical dynamics for small systems. Of course, in the long run, canonical equilibrium values will be recovered by mic.SA because the daemons/walkers’ refresh will vary the total energy.
The performance of mic.SA should also be investigated in problems where the elementary excitation (which is a spin flip in our case) will cause a continuously distributed energy change. The simplest example would be, perhaps, a spin glass with Gaussian-distributed couplings. Probably choosing the daemons/walkers as continuous, exponentially distributed random variables will fully solve the problem, but explicit tests are needed.
Another interesting issue regards the optimal refreshing rate for the daemons/walkers. As we have already explained, if the refresh occurs at a constant time rate, we expect dynamical rescalability with Metropolis. If we refresh at every step with daemons/walkers of infinite capacity, mic.SA is the Metropolis algorithm (if the capacity is limited, the resulting algorithm is just an unusual version of Metropolis). Of course, if the refreshing rate is too low, microcanonical equilibration will occur in between the refreshing steps, which would complicate the dynamical analysis. However, refreshing once every (say) 1000 steps would bring to microcanonical equilibrium only a fairly small system (and the cost of generating random numbers would be divided by a factor of 1000 as compared to a traditional simulation).

### v2

@@ Abstract
Numerical simulations of models and theories that describe complex experimental systems —in fields like high-energy and condensed-matter physics— are becoming increasingly important. Examples include lattice gauge theories, which can describe, among others, quantum chromodynamics (the Standard Model description of strong interactions between elementary particles), and spin-glass systems. Beyond fundamental research, these computational methods also find practical applications, among many others, in optimization, finance, and complex biological problems. However, Monte Carlo simulations, an important subcategory of these methods, are plagued by a major drawback: they are extremely greedy for (pseudo) random numbers. The total fraction of computer time dedicated to random-number generation increases as the hardware grows more sophisticated, and can get prohibitive for special-purpose computing platforms. We propose here a general-purpose microcanonical simulated annealing (MicSA) formalism that dramatically reduces such a burden. The algorithm is fully adapted to a massively parallel computation, as we show in the particularly demanding benchmark of the three-dimensional Ising spin glass. We carry out very stringent numerical tests of the new algorithm by comparing our results, obtained on GPUs, with high-precision standard (i.e., random-number-greedy) simulations performed on the Janus II custom-built supercomputer. In those cases where thermal equilibrium is reachable (i.e., in the paramagnetic phase), both simulations reach compatible values. More significantly, barring short-time corrections, a simple time rescaling suffices to map the MicSA off-equilibrium dynamics onto the results obtained with standard simulations.
@@ I Introduction
Here we propose a widely usable microcanonical simulated annealing (MicSA) formalism, which is both extremely frugal in the number of needed random numbers and fully adapted to a massively parallel computation. The method is demonstrated in the rather demanding case of the three-dimensional Ising spin glass. We carry out our simulations on GPUs to demonstrate that our algorithm nicely parallelizes. We start from a microcanonical ensemble for an extended configuration space that comprises the degrees of freedom we are really interested in (the spins) and some auxiliary degrees of freedom traditionally called daemons. We shall term these auxiliary variables walkers if they are allowed to wander through the system or daemons if they are kept tied to a single lattice site; should the distinction be immaterial, which happens at many points in the discussion, we shall name them daemons/walkers. The reader will find useful analogies between our Monte Carlo formalism and molecular dynamics simulations coupled with a thermostat [30].
@@
Figure 1: After a time rescaling, our microcanonical simulated annealing reproduces canonical dynamics. We study the energy relaxation towards its equilibrium value by plotting the excess energy per spin over its equilibrium value, $e(t)-e_{\infty}$, versus simulation time $t$, as computed for a spin-glass sample on a $160\times 160\times 160$ lattice with binary couplings at temperature $T=0.7\approx 0.64T_{\mathrm{c}}$ ($T_{\mathrm{c}}$ is the critical temperature for the phase transition separating the paramagnetic phase at high temperatures from the spin-glass phase at low temperature). The sample is investigated using two dynamics: a standard, random-number greedy Metropolis algorithm (the label can in the plot stands for canonical; these data were obtained from the Janus II supercomputer) and the new microcanonical simulated annealing (MicSA) algorithm introduced in this work. Specifically, we are using the version of the algorithm with 6 walkers per spin, see Sec. II.8. We plot both data sets, subtracting the same estimate for $e_{\infty}$, which was obtained from the canonical simulation as explained in Appendix A. After the MicSA time is rescaled by the factor $k^{*}=6.034$ (see Table 1), $t^{\text{plot}}=k^{*}\times t^{\text{MicSA}}$, both energy relaxations are indistinguishable (the naif rescaling would have been $k^{\text{naif}}=6$, because of the 6 walkers per spin).
@@
A microcanonical framework with just one daemon was proposed by Creutz [31], precisely to minimize the need for random numbers. Unfortunately, Creutz’s ensemble leads to a rather significant dynamic slowing down in the context of spin glasses [32]. To overcome this dynamic slowdown in non-disordered systems, Rummukainen and other authors proposed using a number of daemons proportional to the number of spins [33, 34, 35, 36]. We shall follow a similar approach, although those authors emphasized equilibrium physics —our focus, instead, is on non-equilibrium dynamics— and they were not particularly concerned with massively parallel simulations. Lustig’s microcanonical ensemble [37] also contains a number of daemons proportional to the number of spins. Lustig’s framework is very effective for studying first-order phase transitions [38], also for disordered systems [39], which suggests a potential effectiveness in the study of spin glasses (see also [33] for a study of first-order transitions in non-disordered systems). This formalism, however, has three major inconveniences: (i) it is unsuitable for parallel computations, (ii) it is not frugal in its usage of random numbers and (iii) it eventually approaches the equilibrium state corresponding to the microcanonical ensemble (which coincides with the canonical ensemble only in the limit of an infinite system size). We cure problems (i) and (ii) by simply skipping the integration of the daemons that Lustig carries out and keeping them instead as explicit dynamic variables. In order to cure problem (iii), we generalize the simulated annealing algorithm [17] to the microcanonical context. We argue (and show by example) that it is sufficient to lower the total energy of the system by acting solely on the daemons at a logarithmic time rate (this is the only part of the algorithm that involves random numbers) to obtain results indistinguishable (after a simple time rescaling) from those produced using a canonical simulation employing the Janus II special-purpose computer [15]. Although rescalability will be considered in much deeper detail below, as a starter, the reader may wish to check Fig. 1.
@@ II.5 Daemons/walkers with a limited capacity
In other words, we should reject any spin-flip that lowers the energy by an amount larger than what the daemon/walker responsible for that update can currently accommodate. On the other hand, having daemons/walkers representable with just two bits reduces the total memory needed. Furthermore, in a multispin coding simulation, the whole algorithm should be implemented employing solely Boolean operations [this includes checking condition Eq. (11) or performing the update in Eq. (9)], which is certainly simpler if the daemons/walkers are representable with only two bits. Refs. [34, 35, 36] also considered daemons with limited capacity and managed them accordingly.
@@ II.6 Daemons or walkers?
Interestingly enough, Refs. [34, 35] implemented their own daemon wandering (although they refreshed their daemons far more often than we do, see Sect. II.7).
@@ II.7 The microcanonical simulated annealing
From the above discussion, it should be clear that refreshing the daemons/walkers configuration at regular times, say every 1000 timesteps (see Sec. II.8 for the precise definition of timestep), should eventually lead any system of finite size to canonical equilibrium. This would dilute the need for random numbers by a factor of 1000, which would probably make the computational weight of generating random numbers irrelevant for most (if not all) applications. However, we shall try to be even thriftier with our random numbers. Indeed, the energy relaxation of spin glasses at and below their critical temperature decreases very slowly with time $t$. The excess energy over the equilibrium value decays as a power law, $H^{\text{\tiny EA}}(t)-\langle H^{\text{\tiny EA}}\rangle_{\beta}\propto N/t^{\epsilon(T)}$, where $\epsilon(T)$ is a fairly small exponent [59, 47] (the exponent is so small that the power law resembles a logarithmic behavior). In fact, for a law $1/\log(t)$, the time required to have a prefixed energy factor decay $f$ grows exponentially with $1/f$. We decided accordingly to refresh our daemons/walkers with At each of those timesteps, the spin system will experience an abrupt, but tiny, energy change as a reaction to the cooler daemons/walkers. This is inappreciable at the scale of Fig. 1, but can be observed with the more refined analysis in Fig. 2-d.
@@ III Results
Figure 2: Comparison of Metropolis and MicSA dynamics with 6 walkers in the spin-glass phase at $T\!=\!0.7\approx 0.64T_{\mathrm{c}}$. a) Time evolution of the difference of the energy at time $t$ and the equilibrium value obtained through a fit to the canonical data. The solid yellow line is for the best fit described in Table 2. b) The coherence length, $\xi_{12}$, as a function of the simulation time for both models. The solid line is again for the best fit described in Table 2. c) Coherence length, $\xi_{12}$, as a function of the ratio between the coherence lengths $\xi_{23}$ and $\xi_{12}$. d) Difference between the results from the two algorithms for the energy, Eq. (22), as a function of the simulation time. e) As in d) but for the coherence length. To compute these differences, we have considered the best fit to our canonical data —solid lines in a) and b)— and the rescaling coefficient $k^{*}=6.034$ (see Appendix A for more information on the estimation of $k^{*}$). Error bars are smaller than data points. For the sake of clarity points are plotted at regular time intervals in logarithmic scale.
@@
Our estimates for $e(t)$, $\xi_{12}(t)$ and the dimensionless ratio $\xi_{23}/\xi_{12}(t)$ are displayed as a function of time in Fig. 2 ($T=0.7\approx 0.64T_{\mathrm{c}}$), Fig. 3 ($T=1.1\approx T_{\mathrm{c}}$) and Fig. 4 ($T=1.4\approx 1.27T_{\mathrm{c}}$). For all three temperatures, we observe that, when $\xi_{12}(t)$ is plotted parametrically as a function of $\xi_{23}/\xi_{12}(t)$, the canonical and the MicSA data fall onto a single curve. This finding indicates that it suffices to consider $\xi_{12}(t)$ when trying to find the correspondence between the $C_{4}(r,t)$ computed with the two dynamics. The figures suggest as well that a shift in $\log t$ (i.e., a simple rescaling for $t$) suffices to superimpose the MicSA data with the canonical curve for both $e(t)$ and $\xi_{12}(t)$. We shall attempt to make this observation quantitatively precise.
@@
Having promoted the canonical $\xi_{12}(t,T)$ to a continuous function, we can simply fit the MicSA curves to
$\xi^{\text{MicSA}}_{12}(t,T)=F^{\xi}_{\text{can}}(k\,t,T)\,,$ (21)
with the rescaling factor $k$ as the only fitting parameter. The best-fit rescaling parameter is named $k^{*}$ in Table 1 and in Figs. 23, and 4. In the figures, we have used for the energy the same value of $k^{*}$ computed with $\xi_{12}$. Each of the different variants of the MicSA algorithm will have its own $k^{*}$ that slightly depends on the temperature. For further details on the computation of $k^{*}$, see Fig. 5 and Appendix A.
Figure 5: Estimation of the best time shift $k^{*}$ at $T_{\mathrm{c}}$. In the main panel, the difference between the logarithm of the original measured time in the MicSA data and the interpolating function $f(k^{*}t)$ as a function of the coherence length $\xi_{12}$. The horizontal solid line at zero indicates the interval of the fit used to estimate $k^{*}$ (see Table 1 for details). Inset: $\log(t)$ as a function of $\xi_{12}$ for both algorithms. Solid lines indicate the best fit for our data. In the case of the MicSA data, the only fitting parameter is $k$, while the rest of the parameters $\Theta=\{a_{0},\,a_{1},\,z\}$ come from the fit to the canonical data described in Appendix A.
@@
$\delta_{e}(t,k)$ $=$ $e^{\text{MicSA}}(t)-F^{e}_{\text{can}}(k\,t,T)\,,$ (22)
$\delta_{\xi}(t,k)$ $=$ $\xi_{12}^{\text{MicSA}}(t)-F^{\xi}_{\text{can}}(k\,t,T)\,,$ (23)
as computed for $k=k^{*}$. These figures show that, when looked at with enough accuracy, $e^{\text{MicSA}}(t)$ is not a smooth function of time (this is most clear for $T\leq T_{\mathrm{c}}$). Indeed, these curves have abrupt (but tiny) changes every time the walkers get refreshed. This is why we have not attempted to get the rescaling factor from the energy. These abrupt changes, as well as the difference $\delta_{e}(t)$ itself, strongly decrease as the time $t$ increases.
@@ IV Conclusions
We have proposed a new Monte Carlo algorithm, microcanonical simulated annealing (MicSA), that is both fully adapted to massively parallel computations and extremely frugal in its use of random numbers. The method is fully general and has a certain number of adaptable features. Chief among these are the number $\gamma$ of daemons/walkers per elementary degree of freedom (spins in our case) and the choice between using daemons or walkers. Walkers have the advantage of avoiding local conservation laws and their resulting slower dynamics, but one can decide whether to use daemons or walkers, and how many, on the view of the application at hand and the hardware platform that will be used.
@@
The method performs fully satisfactorily in the demanding setting of the three-dimensional Ising spin glass, where we have shown that it can reproduce the results of a standard (random-number-greedy) Metropolis algorithm both in and out of equilibrium. In particular, we have shown that the out-of-equilibrium dynamics derived from MicSA can be rescaled onto the one from traditional Metropolis by a simple change in the time unit. The fact that, in the case of walkers, the rescaling factor is nearly equal to the ratio of the number of spin updates in a single timestep for both methods (which coincides with $\gamma$) tells us that a single step of MicSA is worth $\gamma$ Metropolis sweeps, without requiring a single random number.
@@
Our CUDA implementation demonstrates the parallel-friendly nature of the algorithm. Our code, see Appendix B, is publicly available and puts in the hands of a general user a computing performance that could be previously delivered only by the Janus II custom-built computer  [^3] In the context of quantum spin glasses, performance competitive with Janus II has been already obtained on GPUs [25].. Our program may run both standard Metropolis and MicSA, with the MicSA version taking half the time (in fact, in our Metropolis implementation, random-number generation takes about 75% of the simulation time). A standard metric corresponds to the number of picoseconds per spin update. Our MicSA program runs on an H100 GPU at 0.4 ps/spin-update, versus the 3.17 ps/spin-update on a single Janus II FPGA. The above two numbers should, however, be compared with care for two reasons. First, the computing power on the GPU is spread over 128 replicas, while on Janus II updates are concentrated on only two replicas. Hence, for a high-accuracy simulation spanning a limited time range (say up to $t\sim 10^{9}$), the GPU is faster. If, however, one needs to reach times $t\sim 10^{11}$ or larger (which is not uncommon [6]), Janus II still fares better (the GPU might make more spin updates per second by running many replicas, but would need a wall-clock time of many months to finish each simulation). Second, we are, somewhat unfairly, comparing 2025 high-end GPUs with the 2013 Virtex 7 FPGAs that Janus II is made of.
@@
Since MicSA is particularly suitable for high-end FPGAs or even dedicated hardware (based on an ASIC, for instance), we expect that users interested in extreme-scale Monte Carlo simulations will find a significant advantage in our proposal.
@@
The performance of MicSA for system sizes $L$ small enough to be brought to thermal equilibrium also in the spin-glass phase, $T<T_{\mathrm{c}}$, needs to be ascertained. In particular, MicSA needs to be combined with Parallel Tempering [65]. Although we do not foresee difficulties of principle, this algorithmic combination is yet to be tested.
The dynamical properties of MicSA in small systems are likely to be quite peculiar. If the daemon/walkers are refreshed at a very low rate, the system may have enough time to equilibrate in the time lapse between the refreshing steps. Since the equilibrium values of the different observables differ in the canonical and the microcanonical ensemble, it is not obvious that it will always be possible to map MicSA dynamics onto canonical dynamics for small systems. Of course, in the long run, canonical equilibrium values will be recovered by MicSA because the daemons/walkers’ refresh will vary the total energy.
The performance of MicSA should also be investigated in problems where the elementary excitation (which is a spin flip in our case) will cause a continuously distributed energy change. The simplest example would be, perhaps, a spin glass with Gaussian-distributed couplings. Probably choosing the daemons/walkers as continuous, exponentially distributed random variables will fully solve the problem, but explicit tests are needed.
Another interesting issue regards the optimal refreshing rate for the daemons/walkers. As we have already explained, if the refresh occurs at a constant time rate, we expect dynamical rescalability with Metropolis. If we refresh at every step with daemons/walkers of infinite capacity, MicSA is the Metropolis algorithm (if the capacity is limited, the resulting algorithm is just an unusual version of Metropolis). Of course, if the refreshing rate is too low, microcanonical equilibration will occur in between the refreshing steps, which would complicate the dynamical analysis. However, refreshing once every (say) 1000 steps would bring to microcanonical equilibrium only a fairly small system (and the cost of generating random numbers would be divided by a factor of 1000 as compared to a traditional simulation).